OpenStudy (anonymous):

f(x) = 3x2 − 4, what is f(-2)?

Parth (parthkohli):

= (2*-4^2) - 4 = (2*16) - 4 = 32 - 4 = 28

Parth (parthkohli):

Oops, it was 3 * x2, so answer is 48 - 4 = 44

OpenStudy (anonymous):

f(-2)=3*(-2^2)-4 =3*4-4 =12-4 =8

OpenStudy (anonymous):

You really shouldn't try to get medals like that.